# A Million Three-body Binaries Caught by Gaia

### v2

@@ 4 Discussion & Summary
4.1 Comparison to Alternative Binary Formation Channels
Tight field binaries have historically been assumed to form primordially (R. R. King et al., 2012; M. R. Bate, 2014; R. J. Parker & M. R. Meyer, 2014), the remnants of which are the select few binaries that have avoided disruption. Recent state-of-the-art simulations from the Starforge collaboration do in fact find that many binaries form primordially in clusters (D. Guszejnov et al., 2023; J. P. Farias et al., 2024), with a primordial binary fraction ${\sim}40\%$. The majority are bound to the cluster with very wide SMAs and are unlikely to survive the ensuing dynamics in such a dense stellar environment. However, 3BBF may help serve as a complementary formation channel for tight subthermal binaries, the overwhelming majority of which originate almost exclusively in the earliest, densest stage of cluster history within our models. While we do not assert 3BBF as a prominent method for tight binary generation, primordial binary formation need not be the only means.
Wide Gaia binary formation channels have been more elusive, primarily because the proposed formation channels may produce many wide thermal binaries—e.g., cluster dissolution (M. B. N. Kouwenhoven et al., 2010; N. Moeckel & M. R. Bate, 2010; J. Peñarrubia, 2021)—or a few superthermal binaries—e.g., disk fragmentation (J.-E. Lee et al., 2017; S. Xu et al., 2023)—but not both in the required quantities. As highlighted in § 3.2, the tendency for binary field evolution is to accelerate the widening of binaries towards disruption. This implies that any prospective wide binary formation channel matching the observed Gaia distribution before applying field evolution cannot be the dominant source of the Gaia wide binary population. Such is the case for binary formation by cluster dissolution, preferentially forming extremely wide binaries before considering field evolution.
With 3BBF, the preferred outcome is the formation of highly eccentric wide binaries. Accounting for the disruptive capacity in clusters and field perturbations, 3BBF can circumvent many of the formation challenges, implying that 3BBs may contribute a significant fraction of wide binaries observable today.
@@ 4.2 Limitations and Future Work
For a first foray into assembling a Galactic 3BBF rate, we opted for a simple, physically motivated astrophysical model. Yet, there are several practical developments that could increase the model’s physical realism and self-consistency.
@@
While we have devoted much effort to modeling binary destruction in the birth cluster (§ A.5) and Galactic field (§§ C.1.1 and C.2), we do not consider the effect of in-cluster dynamical interactions not fully disrupting an escaping 3BB en route out of its natal cluster. As with field evolution, these dynamical interactions will typically widen binaries that are as low-mass and wide as the ones we consider, while further thermalizing their eccentricities (D. C. Heggie, 1975). However, our conclusions are unchanged because additional thermalization can only push the 3BB eccentricity distributions in Figure 5 towards the more thermal Gaia eccentricities identified by H.-C. Hwang et al. (2022b). Additionally, while we incorporate a simple giant molecular cloud disruption recipe (§ C.2), we do not consider the possibility of disruption from other MW substructures, like dark matter subhalos or other star clusters. Incorporating said structures alongside a time-dependent density profile of the MW would help improve the physical accuracy of our framework and increase the destruction rate of 3BBs in the Galactic field.
@@
Our methodology also neglects tidal circularization and orbital decay from isolated binary stellar evolution (BSE), such as common envelope physics. BSE would be especially important modeling the large concentration of extremely eccentric binaries expected to be produced by 3BBF (J. Stegmann et al., 2024). Predicting the stellar metallically of binary members would also require an accurate treatment of BSE, which we leave for future work.
@@
Combining the cluster evolution equations from M. Gieles et al. (2011) and an equal-mass Plummer model is a powerful and convenient way to rapidly characterize and evaluate cluster properties across mass scales purely analytically. However, a more refined prescription for the structure and evolution of star clusters with realistic initial mass functions and stellar populations could greatly improve the physical realism of our model. Notably, black holes (BHs) play critical roles in the evolution of star clusters (D. Merritt et al., 2004; A. D. Mackey et al., 2007, 2008; P. G. Breen & D. C. Heggie, 2013; M. Morscher et al., 2015; M. Peuten et al., 2016; S. Chatterjee et al., 2017; K. Kremer et al., 2019; N. C. Weatherford et al., 2020, 2021; M. Gieles et al., 2021; M. Gieles & O. Y. Gnedin, 2023; D. Roberts et al., 2025). Their high masses, low velocities, and high local density after segregating to the cluster center make BHs the primary catalyst for assembling compact object 3BBs (e.g., N. Ivanova et al., 2005; M. Morscher et al., 2015; N. C. Weatherford et al., 2023)—and possibly for wider 3BBF more generally.
@@
Modeling 3BBF in multi-mass clusters will require a semi-analytic treatment of unequal mass 3BBF. Exploring unequal-mass 3BBF for the first time, D. Atallah et al. (2024) determined that the pairing probabilities between unequal-mass bodies in a positive energy encounter generally favor the pairing of the two least massive bodies. This revised the prior notion that the two most massive bodies are the most likely to bind (e.g., M. Morscher et al., 2015), which is the case for three-body encounters with negative total energy, such as binary–single interactions (N. C. Stone & N. W. C. Leigh, 2019; Y. B. Ginat & H. B. Perets, 2021; B. Kol, 2021). Unequal-mass 3BBF also features subtle changes to the overall SMA and eccentricity distributions of 3BBs. Yet, the critical characteristics imparted by 3BBF, that (i) wide binaries tend to be superthermal and (ii) binaries with an SMA smaller than the radius of the interaction volume tend to be described by a thermal distribution, is not changed.
The degree to which these distributions diverge from equal-mass 3BBF is small for mass ratio variance within a factor of ${\sim}2$, so a significant modification to the overall demographic of Gaia binaries is unexpected. Hard binaries pairing a compact object with a star will be most affected by a proper unequal-mass treatment of 3BBF, as these systems tend to exhibit mass ratios far from unity. Unfortunately, the parameter space encompassing unequal-mass 3BBF is simultaneously immense and subtle. A semi-analytic fit appropriate for our G3R framework, or any other population-wide study, requires a dedicated investigation. Full $N$-body simulations will be important to mitigate uncertainties in employing the current semi-analytic model with low-mass clusters; in future work, we plan to conduct such simulations.
@@
Our models are likely insufficient at characterizing binary formation during cluster dissolution, whether through 3BBF or more complicated scattering modes. M. B. N. Kouwenhoven et al. (2010) estimated a wide binary fraction present at cluster dissolution of $10\%$ to $20\%$. This corresponds to one to two wide binaries formed per cluster at dissolution, though N. Moeckel & C. J. Clarke (2011) found integrating their models forward in time by a factor of two diminished the binary survival rate by an order of magnitude. The overwhelming majority of these binaries would likely be destroyed in the field (with typical initial SMA $>5000\,\rm{AU}$), but the surviving few would boost the overall wide binary formation rate and enhance the relative contribution from short-lived low-mass clusters.
@@ 4.3 Summary
We have developed a semi-analytic Galactic three-body binary formation rate (G3R) and demonstrate that three-body binary formation (3BBF) efficiently produces highly eccentric wide binaries. Our model predicts that binaries with SMA ${<}10^{3}\rm AU$ exhibit subthermal eccentricities whereas binaries with SMA ${>}10^{3}\rm\,AU$ are superthermal. Without fine-tuning, this framework closely matches the observed SMA distribution of Gaia’s wide binaries (K. El-Badry et al., 2021) and qualitatively matches the eccentricity distributions of H.-C. Hwang et al. (2022b) when applying cumulative scatter binary field evolution constructed by C. Hamilton & S. Modak (2024). The predicted number of wide binaries in the solar neighborhood matches or exceeds that observed by Gaia, though we have not accounted for observational incompleteness or selection effects. However, these systematics primarily affect Gaia’s ability to observe binaries with separations $<200\rm AU$. Thus, 3BBF may be an essential formation channel for the widest field binaries.
@@
We are grateful for insightful conversations shared with J. Binney, N. Choski, N. Frankel, M. Gieles, C. Johnson, B. Kocsis, K. Kremer, J. Li, A. Marszewski, J. Magorrian, C. O’Connor, T. Panamarev, H.B. Perets, S. Rose, M. Rozner, B. Scott, and N.C. Stone on topics ranging from cluster formation history to Milky Way evolution and dynamics. D.A. would like to thank Fred Rasio for his commitment and support towards the production of this work and the author’s doctorate. D.A. acknowledges support from the CIERA Board of Visitors Fellowship. The work of Y.B.G. was supported in part by a Leverhulme Trust International Professorship Grant to S.L. Sondhi (No. LIP-2020-014), and partly by a Simons Investigator Award to A.A. Schekochihin. This work used computing resources at CIERA funded by NSF PHY-2406802. This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
@@ A.1 The 3BBF Rate
From D. Atallah et al. (2024), the volumetric rate of encounters between three unbound bodies (3UBs) for a field of equal masses $m$ with a Maxwellian velocity distribution is
@@
The 3BBF rate is simply the 3UB encounter rate multiplied by the numerically determined probability of binary formation, $P_{F}$, in 3UB encounters with the given $n$, $R_{1}$, and $\sigma$. D. Atallah et al. (2024) calculated this semi-analytic 3BBF rate with similar initial scattering conditions to S. J. Aarseth & D. C. Heggie (1976a), setting $R_{2}=15R_{1}$. Here, $R_{2}$ is the offset distance between the slowest moving body and the interaction volume; the other two bodies are initiated further from $R_{1}$, at distances proportional to their randomly drawn initial velocity (assuming a Maxwellian velocity distribution with dispersion $\sigma$). [^7] See D. Atallah et al. (2024) for a more thorough numerical exploration of the simultaneous scattering of three unbound bodies (3UB). In a dense cluster environment, however, if $R_{1}\sim r_{\rm sep}$, the average local interparticle distance, then it does not make sense for bodies to travel a distance $R_{2}=15R_{1}$; the individual bodies would likely have a strong encounter en route to the 3UB interaction, breaking the approximation of an isolated encounter. The equal-mass 3BBF scattering probability of D. Atallah et al. (2024) thus needs to be refined by reducing $R_{2}/R_{1}$ from $15$ to $3$. The resulting numerically determined 3BBF probability is
$P_{F}$ $=c_{0}\left(\frac{2b_{90}}{R_{1}}\right)^{2}=c_{0}\left(\frac{2Gm}{3\sigma^{2}R_{1}}\right)^{2},\ c_{0}=3.82,$ (A2)
@@
For convenience, $R_{1}$ is rewritten in terms of the average interparticle separation, $r_{\rm sep}=\left(\frac{3}{4\pi n}\right)^{1/3}$, becoming $R_{1}=kr_{\rm sep}$. The $G^{2}m^{2}n^{2}\sigma^{-3}$ scaling above is notably shallower than the classic $G^{5}m^{5}n^{3}\sigma^{-9}$ scaling found in detailed balance calculations (D. C. Heggie, 1975; J. Goodman & P. Hut, 1993) and recovered recently in our modern approaches (Y. B. Ginat & H. B. Perets, 2024; D. Atallah et al., 2024), which applies to the formation of hard binaries. Prior estimates fixate on hard 3BBF, whereas we enable both locally hard and soft 3BBF. By relaxing the limit on binding energy and, instead, enforcing a geometric limit based on particle separations, we recover a formation rate scaling far more enabling of 3BBF, albeit with arbitrarily small binding energies. In the limit of hard binary formation, the combination of Equations (A3) & (A7, see § A.2) yields the classic scaling.
@@
[…]

### v3

@@ 4 Discussion & Summary
4.1 Comparison to Alternative Binary-Formation Channels
Tight field binaries have been historically assumed to form primordially (R. R. King et al., 2012; M. R. Bate, 2014; R. J. Parker & M. R. Meyer, 2014), the remnants of which are the select few binaries that have avoided disruption. Recent state-of-the-art simulations from the Starforge collaboration do in fact find that many binaries form primordially in clusters (D. Guszejnov et al., 2023; J. P. Farias et al., 2024), with a primordial binary fraction of ${\sim}40\%$. The majority are bound to the cluster with very wide SMAs and are unlikely to survive the ensuing dynamics in such a dense stellar environment. However, 3BBF may help serve as a complementary formation channel for tight subthermal binaries, the overwhelming majority of which originate almost exclusively in the earliest, densest stage of cluster history within our models. While we do not assert 3BBF as a prominent method for tight binary generation, primordial binary formation need not be the only means.
Wide Gaia binary formation channels have been more elusive, primarily because the proposed formation channels readily produce many wide binaries—e.g., cluster dissolution or random pairing (M. B. N. Kouwenhoven et al., 2010; N. Moeckel & M. R. Bate, 2010; J. Peñarrubia, 2021)—or a few highly eccentric binaries—e.g., disk fragmentation (J.-E. Lee et al., 2017; S. Xu et al., 2023)—but not necessarily both in the required quantities. In particular, before this work, there had not yet been a thorough investigation of the cluster formation and dissolution channel’s contribution to the present-day field binary population. As highlighted in § 3.2, the tendency for binary field evolution is to accelerate the widening of binaries towards disruption. This implies that any prospective wide binary formation channel matching the observed Gaia distribution, prior to the application of field evolution, cannot be the dominant source of the Gaia wide binary population.
With 3BBF, the formation of highly eccentric wide binaries is preferred, especially late in a cluster’s life—supporting the cluster dissolution channel. By accounting for the disruptive capacity of dynamical encounters in clusters and galactic field perturbations, we find that 3BBF may contribute a significant fraction of Galactic wide binaries observable today.
@@ 4.2 Limitations and Future Work
For a first foray into assembling a Galactic 3BBF rate, we opted for a simple, physically motivated astrophysical model. That said, there are several practical developments that can further improve the physical realism and self-consistency of our method.
@@
Substantial effort is already devoted to modeling binary destruction in the natal cluster (§ A.5) and in the Galactic field (§§ C.1.1 and C.2). What we do not follow is the effect of in-cluster dynamical interactions not fully disrupting a 3BB, instead opting to forcefully discount any binary that undergoing even a weak encounter. These dynamical interactions will typically widen low-mass, wide binaries and further thermalize their eccentricities (D. C. Heggie, 1975). Ignoring them does not alter our conclusions: the few wide binaries not disrupted in-cluster would provide a subtle thermalization that would only push the 3BB eccentricity distributions in Figure 4 towards the more thermal Gaia eccentricities identified by H.-C. Hwang et al. (2022b). Additionally, while we incorporate a simple giant molecular cloud disruption recipe (§ C.2), we do not consider the possibility of disruption from other MW substructures, like dark matter subhalos or other star clusters. Incorporating said structures alongside a time-dependent density profile of the MW would help improve the physical accuracy of our framework and increase the destruction rate of 3BBs in the Galactic field.
@@
Tidal circularization, orbital decay, and common-envelope episodes from isolated binary stellar evolution (BSE), has also been neglected. BSE is especially important for the extremely eccentric binaries readily produced by 3BBF (J. Stegmann et al., 2024). Predicting the stellar metallically of binary members also requires an accurate treatment of BSE, which we leave for future work.
@@
Combining the cluster evolution equations from M. Gieles et al. (2011) and an equal-mass Plummer model is a fast and transparent means to rapidly evaluate cluster properties across mass scales. However, a more refined prescription for the structure and evolution of star clusters with realistic initial mass functions and stellar populations will greatly improve the physical realism of our model. Notably, black holes (BHs) play critical roles in the evolution of star clusters (D. Merritt et al., 2004; A. D. Mackey et al., 2007, 2008; P. G. Breen & D. C. Heggie, 2013; M. Morscher et al., 2015; M. Peuten et al., 2016; S. Chatterjee et al., 2017; K. Kremer et al., 2019; N. C. Weatherford et al., 2020, 2021; M. Gieles et al., 2021; M. Gieles & O. Y. Gnedin, 2023; D. Roberts et al., 2025). After segregating to the cluster core, their high masses, low velocities, and high local density make BHs the primary catalyst for assembling compact object 3BBs (e.g., N. Ivanova et al., 2005; M. Morscher et al., 2015; N. C. Weatherford et al., 2023)—and possibly for wider 3BBF more generally.
It is unlikely compact objects comprise many of the wide, eccentric binaries which are the focus of this work. As discussed in § 3.2, wide superthermal 3BBs that survive and successfully migrate to the solar neighborhood overwhelmingly originate outside of the core of low-mass clusters $(M_{\rm cl0}\lesssim 10^{3}M_{\odot})$ and are catalyzed within the final ${\sim}20\%$ of the cluster’s life. These clusters would host only a few compact objects, which primarily reside in the cluster core. We therefore expect the surviving wide 3BBs to be mostly low-mass stellar pairs. Additionally, the interaction cross-section of wide binaries born outside the cluster’s core is dominated by single stars, encounters our framework already accounts for conservatively (see § A.5). Thus, the key benefit to incorporating a richer stellar population would be to uncover subtle pairing tendencies between different stellar types.
@@
Modeling 3BBF in multi-mass clusters will require a semi-analytic treatment of unequal mass 3BBF. The first such study by D. Atallah et al. (2024) determined that positive energy 3UB encounters generally favor the pairing of the two least massive bodies. This revised the prior notion that the two most massive bodies are the most likely to bind (e.g., M. Morscher et al., 2015), which is the case for three-body encounters with negative total energy, such as binary–single interactions (N. C. Stone & N. W. C. Leigh, 2019; Y. B. Ginat & H. B. Perets, 2021; B. Kol, 2021). Unequal-mass 3BBF produces only modest shifts in SMA and eccentricity distributions when mass ratios vary by ${\lesssim}2$, so the overall Gaia-binary demographics should not change dramatically. The largest impact should appear in hard, compact-object–plus-star binaries, whose extreme mass ratios demand an explicit unequal-mass treatment. The required parameter space is vast and subtle, calling for a dedicated study and, ideally, full cluster $N$-body simulations—an effort we plan to undertake.
@@
Our models are likely insufficient at characterizing binary formation late into cluster dissolution, whether through 3BBF or more complicated scattering modes. M. B. N. Kouwenhoven et al. (2010) estimated a wide binary fraction present at cluster dissolution of $10\%$ to $20\%$. This corresponds to one to two wide binaries formed per cluster at dissolution, though N. Moeckel & C. J. Clarke (2011) found integrating their models forward in time by a factor of two diminished the binary survival rate by an order of magnitude, leaving only a fraction of clusters emitting a single binary at dissolution. This finding is consistent with our own despite approximating cluster evolution with a simplistic analytic model. Whether the binaries form in the final moments of the cluster’s life, or somewhat earlier, through isolated 3BBF is left for a future investigation.
Finally, our model already closely matches or exceeds Gaia’s DR3 binary counts despite applying several conservative filters for binary survival. More complicated channels tied to cluster dissolution (e.g., M. B. N. Kouwenhoven et al. (2010); N. Moeckel & C. J. Clarke (2011)) will overlap with 3BBF because realistic dissolution makes no such distinction between dynamical formation mechanisms. Any concern surrounding over-counting should be allayed by the eventual inclusion of multi-mass 3UB scattering. The 3BBF rate decreases as the mass ratio between the most- and least-massive scattering bodies diverges from unity (D. Atallah et al., 2024), favoring 3BBF with comparable masses, thus reducing the overall 3BBF rate. Additionally, the current Gaia binary counts are likely an underestimate (K. El-Badry & H.-W. Rix, 2018; K. El-Badry et al., 2021), so we expect the catalog of wide binaries to expand in Gaia DR4, potentially pushing Gaia observations and our theoretical results into greater agreement.
@@ 4.3 Summary
By developing a semi-analytic Galactic three-body binary formation rate (G3R), we have demonstrated that three-body binary formation (3BBF) can efficiently generate highly eccentric, wide field binaries. Our model predicts that binaries with SMA $10^{3}\rm AU$ exhibit subthermal eccentricities whereas binaries with SMA $10^{3}\rm\,AU$ are superthermal. Contingent on physically-motivated assumptions regarding cluster-structure and tides, this framework closely matches the observed SMA distribution of Gaia’s wide binaries (K. El-Badry et al., 2021) and the eccentricity distributions of H.-C. Hwang et al. (2022b) when applying dynamical binary field evolution framework constructed by C. Hamilton & S. Modak (2024). The predicted number of wide binaries in the solar neighborhood matches or exceeds that observed by Gaia, though we have not accounted for observational incompleteness or selection effects. However, these systematics primarily affect Gaia’s ability to observe binaries with separations $200\rm AU$. Thus, 3BBF may be an essential formation channel for the widest field binaries.
@@
We are grateful for insightful conversations shared with J. Binney, N. Choksi, N. Frankel, M. Gieles, C. Johnson, B. Kocsis, K. Kremer, J. Li, A. Marszewski, J. Magorrian, C. O’Connor, T. Panamarev, H.B. Perets, S. Rose, M. Rozner, B. Scott, and N.C. Stone on topics ranging from cluster formation history to Milky Way evolution and dynamics. D.A. would like to thank Fred Rasio for his commitment and support towards the production of this work and the author’s doctorate. D.A. acknowledges support from the CIERA Board of Visitors Fellowship. This work was supported in part by a Leverhulme Trust International Professorship Grant (No. LIP-2020-014). The work of Y.B.G. was partly supported by a Simons Investigator Award to $A.$ Schekochihin. This work used computing resources at CIERA funded by NSF PHY-2406802. This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
@@ A.1 The 3BBF Rate
From D. Atallah et al. (2024), the volumetric rate of encounters between three unbound bodies (3UBs) for a field of equal masses, $m$, with a Maxwellian velocity distribution is
@@
The 3BBF rate is simply the 3UB encounter rate multiplied by the numerically determined probability of binary formation, $P_{F}$, in 3UB encounters with the given $n$, $R_{1}$, and $\sigma$. D. Atallah et al. (2024) calculated this semi-analytic 3BBF rate with similar initial scattering conditions to S. J. Aarseth & D. C. Heggie (1976), setting $R_{2}=15R_{1}$. Here, $R_{2}$ is the offset distance between the slowest moving body and the interaction volume. The other two bodies are initiated further from $R_{1}$, at distances proportional to their randomly drawn initial velocity, assuming a Maxwellian velocity distribution with dispersion $\sigma$. [^6] See D. Atallah et al. (2024) for a more thorough numerical exploration of the simultaneous scattering of three unbound bodies (3UB). In a dense cluster environment, if $R_{1}\sim r_{\rm sep}$, the average local interparticle distance, then it does not make sense for bodies to travel a distance $R_{2}=15R_{1}$. The individual bodies would likely have a strong encounter en route to the 3UB interaction, breaking the approximation of an isolated encounter. The equal-mass 3BBF scattering probability of D. Atallah et al. (2024) thus needs to be refined by reducing $R_{2}/R_{1}$ from $15$ to $3$. The resulting numerically determined 3BBF probability is
@@
For convenience, $R_{1}$ is rewritten in terms of the average interparticle separation, $r_{\rm sep}=\left(4\pi n/3\right)^{-1/3}$, becoming $R_{1}=k\,r_{\rm sep}$. The $G^{2}m^{2}n^{2}\sigma^{-3}$ scaling above is notably shallower than the classic $G^{5}m^{5}n^{3}\sigma^{-9}$ scaling found in detailed balance calculations (D. C. Heggie, 1975; J. Goodman & P. Hut, 1993) and recovered recently in our modern approaches (Y. B. Ginat & H. B. Perets, 2024; D. Atallah et al., 2024), which applies to the formation of hard binaries. Prior estimates fixate on hard 3BBF, whereas we enable both locally hard and soft 3BBF. By relaxing the limit on binding energy and, instead, enforcing a geometric limit based on particle separations, we recover a formation rate scaling far more enabling of 3BBF, albeit with arbitrarily small binding energies. In the limit of hard binary formation, the combination of Equations (A3) & (A7, see § A.2) yields the classic scaling.
@@
It is straightforward to integrate over all $k$, but, realistically, tidal forces from the cluster will tear apart an encounter before it occurs if $k$ is large enough. To account for this, a transformation can be applied to values of $k$ appearing in Equation (A3), limiting the maximum possible interaction volume self-consistently. A convenient choice for the maximum value of $k$ is some fraction, $\eta_{\rm H}$, of the Hill Radius, $R_{H}\approx\,r\left(\langle m\rangle/M_{\rm cl,enc}\right)^{1/3}$; $M_{\rm cl,enc}$ is the total cluster mass enclosed at the radial location of the encounter. In terms of $r_{\rm sep}$, the fraction of local Hill Radius in a Plummer cluster is $k_{H}=\eta_{H}(1+(r/b)^{2})^{-1/3}$, where $\eta_{H}$ is arbitrarily selected to be $\ll 1$; we explore the effects of conservatively limiting $\eta_{H}$ in our results (§ 3). The transformation applied is thus: $k\rightarrow k_{1}=\min(k,k_{H})$. Specified to the case of encounters within a single field, and including the Hill Radius limit, Equation (A3) becomes [^7] Applying the $k_{1}$ transformation is equivalent to evaluating the integral $\int_{0}^{k}\tilde{\Gamma}_{F,0}(k_{1})P_{\rm iso}(k)dk=\int_{0}^{k_{H}}\tilde{\Gamma}_{F,0}(k)P_{\rm iso}(k)dk+\tilde{\Gamma}_{F,0}(k_{H})\int_{k_{H}}^{k}P_{\rm iso}(k)dk$. The transformation should not be applied to Equation (A5) since it is only the distribution of all possible isolated encounters and should integrate to 1 over all $k$-space independent of the local physics in an environment. In the limit that $k_{H}^{3}\ll 1$, the integral reduces to $\int_{0}^{k}\tilde{\Gamma}_{F,0}(k_{1})P_{\rm iso}(k)dk\approx\tilde{\Gamma}_{F,0}(k_{H})$.